OpenStudy (anonymous):

Which fractions are equivalent to -4/9.

TheSmartOne (thesmartone):

\(\sf\Large-\frac{4}{9}\times\frac{x}{x}=?\) Put in any value of x, it will always be equivalent to -4/9

TheSmartOne (thesmartone):

Put in any value for x***